# TexHOI: Reconstructing Textures of 3D Unknown Objects in Monocular Hand-Object Interaction Scenes

### v1

@@ 4.1.4 Pose Refinement
where $C(r)$ and $\hat{C}(r)$ are the predicted and groundtruth RGB colors along the ray $r$. Hand, object, and background masks are ensured consistent with ground truth by calculating the loss between predicted segmentation one-hot label $S(r)$, calculated using Eq. 8, and estimated ground truth segmentation [cheng2023segment] $\hat{S}(r)$:
$\mathcal{L}_{seg1}=\Sigma_{r}||S(r)-\hat{S}(r)||.$ (10)
@@
where all $\lambda_{*}$ are weights of the losses empirically assigned.
@@
Figure 2: After Stage 1, for each (a) input image, predicted (b) object mask and (c) hand-object mask are calculated.
For a given frame (Fig. 2 (a)), this stage prepares the object’s implicit network, hand and object pose, object mask (Fig. 2 (b)) and hand-object mask (Fig. 2 (c)) for the second stage of our method.
@@ 4.2.1 Environment Illumination
$L_{d}(\omega_{i})=\Sigma_{j=1}^{128}G(\omega_{i};\psi_{j},\mu_{j},\lambda_{j}).$ (13)
@@ 4.2.3 Hand Occlusion
One of the key challenges in rendering hand-object interactions is handling occlusion caused by the hand. To model this, we represent the hand as a set of $108$ parameterized spheres similar to power cells [wang2022computing]. We first place all spheres $\{c_{i},r_{i}\}_{i=1}^{108}$ manually in the canonical MANO hand volume $\Psi$, with $c_{i}$ and $r_{i}$ being their centers and radii. The sphere centers $\{c_{i}\}_{i=1}^{108}$ are determined on the hand skeleton, manually selected between the MANO joints. This allows the spheres to maintain the hand structure whenever the pose parameter of the MANO hand model is deformed. The sphere radii $\{r_{i}\}_{i=1}^{108}$ requires partitioning the surface of the hand model into individual spheres. These spheres define a power diagram that partitions the MANO hand volume $\Psi$ into a set of power cells. Each power cell $\Psi_{i}^{pow}$ consists of the points $x$ that are closest to a particular sphere $(c_{i},r_{i})$:
@@
where $d_{pow}(x,c_{i},r_{i})=||x-c_{i}||^{2}-r_{i}^{2}$ is the power distance between the point $x$ and the sphere $(c_{i},r_{i})$. The power diagram partitions all vertices $\{v_{j}\}$ of canonical MANO hand model into each individual set $\mathcal{V}_{i}=\{v_{j}|v_{j}\in\Psi_{i}^{pow}\}$. Whenever the MANO hand model is deformed by its pose and shape parameters, the center $\hat{c}_{i}$ and radius $\hat{r}_{i}$ of each sphere can be updated correspondingly by the following rule:
$\begin{split}\hat{c}_{i}&=\frac{1}{|\mathcal{V}_{i}|}\sum_{v_{j}\in\mathcal{V}_{i}}\hat{v}_{j},\\ \hat{r}_{i}&=\frac{1}{|\mathcal{V}_{i}|}\sum_{v_{j}\in\mathcal{V}_{i}}||\hat{n}_{j}\cdot(\hat{v}_{j}-\hat{c}_{i})||,\end{split}$ (17)
@@
This granular representation allows us to model how different parts of the hand obstruct the light and cast shadows on the object’s surface. For each surface point on the object, we divide its lobe of Spherical Gaussian (SG) into a grid of 64x64 patches, each representing a small region of the environment illumination and BRDF. Each one of the $108$ spheres representing MANO hand is projected onto the SG lobe, and the patches belonging to the occluded regions are collected. Finally, the fractional value of the integral of the occluded patches gives a measure of the extent to which each part of the hand blocks light from reaching the object. For the following integral of Spherical Gaussian [zhang2021physg] with lobe sharpness $\lambda$,
@@
$S(\theta,\phi,\lambda)$ represents the accumulated Spherical Gaussian over a patch of the hemisphere, where $\theta$ and $\phi$ are angular coordinates. This means when $\theta=\pi$ and $\phi=\pi$, it covers the whole hemisphere. To compute the fractional value for a given angular region, we normalize by dividing the integral result by $S(\pi,\pi,\lambda)$ [iwasaki2012real], as shown below:
$\hat{S}(\theta,\phi,\lambda)=\frac{S(\theta,\phi,\lambda)}{S(\pi,\pi,\lambda)}.$ (19)
This normalized value helps us calculate the proportion of the spherical Gaussian that falls within a specific patch of the sphere. Now, for a patch defined by the angular coordinates $(\theta_{0},\phi_{0})$, $(\theta_{1},\phi_{0})$, $(\theta_{0},\phi_{1})$, and $(\theta_{1},\phi_{1})$ (See Fig. 4), the final output is determined by computing the contributions from the corners of the patch:
$\begin{split}O(x)=\hat{S}(\theta_{1},\phi_{1},\lambda)-\hat{S}(\theta_{1},\phi_{0},\lambda)-\hat{S}(\theta_{0},\phi_{1},\lambda)\\ +\hat{S}(\theta_{0},\phi_{0},\lambda).\end{split}$ (20)
This formula computes the difference between the normalized Spherical Gaussian values at the corners of the patch, giving us the contribution of the patch itself. Therefore, $O(x)$ at the surface point $x$ gives the fractional value of the patch’s contribution in occlusion computation as shown in Fig. 5.
@@
This approach, following Iwasaki et. al. [iwasaki2012real] not only mitigates the computational overhead associated with ray tracing but also ensures that the occlusion effects are incorporated into the final rendering with minimal loss of realism. Moreover, since it is directly computed over the Spherical Gaussian (SG) lobe at $x$, $O(x)$ can be directly used as a constant term with other SG terms.
@@ 4.2.4 Final Illumination
We discussed direct illumination $L_{d}(\omega_{i})$ from environment, along direction $w_{i}$ in Sec. 4.2.1, and hand occlusion $O(x)$ in Sec. 4.2.3. In the regions of hand occlusion, the hand casts its own indirect illumination, which can be perceived as the hand’s reflection on the object’s surface. Hence, we learn a constant indirect illumination $L_{i}$, starting from a generic skin color (#$\#E0AC69$) from an online search. The final illumination $L(\omega_{i},x)$ is thus obtained as follows:
$L(\omega_{i},x)=L_{d}(\omega_{i})*(1-O(x))+L_{i}*O(x).$ (21)
@@
Combining illumination $L(\omega_{i},x)$ from Eq. (21), BRDF $\phi(x;\omega_{o},\omega_{i})$ from Eq. (14), and cosine term $\omega_{i}\cdot n$ from Eq. (22), we get the final color $c(\omega_{o},x)$ at the surface point $x$ with view direction $\omega_{o}$ as the integral equation given in PBR Eq. (1). To maintain the consistency of object shape, and avoid the hand’s involvement in object geometry and texture reconstruction, we propose the following loss functions. An RGB loss is calculated on the color $c^{ho}$ predicted on $N_{ho}$ sampled pixels in the intersection of the predicted object mask and groundtruth hand-object mask (Fig. 2):
@@
where $\hat{c}_{i}$ is groundtruth color of the $N_{ho}$ sampled pixels. The below mask loss calculates minimum SDF $S^{no}$ on the $N_{no}$ rays corresponding to the sampled pixels outside the intersection of the predicted object mask and ground truth object mask (Fig. 2):
@@
Just like in stage 1, $\mathcal{L}_{eikonal2}$ regularizes object’s geometry. The final loss optimized , with $\lambda$ values empirically assigned is as follows:
$\mathcal{L}_{s2}=\mathcal{L}_{rgb2}+\lambda_{mask2}\mathcal{L}_{mask2}+\lambda_{eikonal2}\mathcal{L}_{eikonal2}.$ (25)
@@ 4.3 Implementation Details
We utilize monocular camera capture datasets that encompass all viewpoints around the object, ensuring comprehensive coverage for high-fidelity texture and geometry learning. For the first stage, the 3D annotations estimated include hand-object pixel-wise segmentation [cheng2023segment, kirillov2023segment, yang2022deaot, yang2021aot, liu2023grounding, gong21b_interspeech], hand MANO pose estimation [lin2021end], and object pose estimation [sarlin2019coarse, sarlin2020superglue] in the camera-coordinate system.
The first stage of our pipeline involves the composite volumetric rendering of the hand, object, and background, following a pipeline similar to HOLD [fan2024hold], using the initial estimates of the hand and object poses, along with segmentation masks. HOLD optimizes hand and object poses, geometry, and an initial estimate of texture.
Using the optimized poses and object geometry from Stage 1, we derive the object mask and hand-object segmentation mask. These are then used in the second stage, which focuses on physics-based surface rendering of the object, following a pipeline inspired by PhySG [zhang2021physg]. A key modification in our method is the computation of occlusion to account for hand-object interaction, based on the approach in Iwasaki et al. [iwasaki2012real]. Prior works, such as PhySG, do not compute occlusion in this way, setting our method apart. In Stage 2, we optimize the following key parameters at a surface point $x$:
@@
Figure 7: Qualitative Comparison: TexHOI albedo predictions avoid baked hand-occlusions and environment-illumination. Hand occlusions can be observed in “Hand-Object Interaction”. The final TexHOI RGB is calculated using Physics-Based Rendering, and is compared with other state-of-the-art approaches, HOLD [fan2024hold] and InvRender [zhang2022modeling].
@@
Table 1: Quantitative comparison is performed between our method, TexHOI, and other state-of-the-art approaches that reconstruct geometry and texture. HOLD bakes hand-occlusions and illumination on its texture prediction, InvRender does not handle dynamic occlusions.
Figure 8: Ablation studies compare our model with the versions trained without indirect illumination, and without any occlusions. Illumination effects on surface albedo, and hand occlusions on surface can be observed.
@@
Table 2: Quantitative ablative comparison of our model with the versions without indirect illumination, and without hand occlusions. Without illumination, the interacting hand casts a darker texture. Without occlusion handling, artifacts occur in major hand grasp regions.
@@ 5.1 Dataset
  • HO3D: The training data of HO3D [hampali2020honnotate] dataset contains 9 objects from YCB [xiang2017posecnn] models. We use monocular frames of six different objects from the dataset. These six objects were specifically selected based on the presence of sufficient texture details, which is essential for evaluating the effectiveness of our texture reconstruction. The six objects (and their subjects) are - sugar box (“ShSu”), cracker box (“MC”), bleach cleanser (“ABF”), meat can (“GPMF”), mustard bottle (“SM”), and drilling machine (“MDF”). The $3$ left objects in training data - scissors, banana, and mug contain no detailed-texture information for comparison.
  • In-the-wild: We also use monocular frame sequences of three objects captured in real-world settings (“in-the-wild”) [fan2024hold]. These objects provide further diversity in testing by presenting challenging hand-object interactions and varied environmental conditions. The objects are - bottle, kettle, and rubrics cube.
@@ 5.2 Baselines
We use HOLD [fan2024hold] as our primary baseline for performance comparison due to its proven effectiveness in dynamic hand-object interaction tasks, where it outperforms methods such as DiffHOI [ye2023diffusion] and BundleSDF [wen2023bundlesdf]. This makes HOLD an appropriate benchmark for assessing the improvements introduced by our method. Our approach, however, goes beyond what HOLD achieves by focusing on texture reconstruction and illumination effects, areas that HOLD does not fully address. To highlight the improvements, we compare our full pipeline results with HOLD alone. This comparison demonstrates how our method enhances texture accuracy, particularly in capturing the complex interactions between hands and objects, such as shadows, reflections, and visibility changes, which are often overlooked in prior works. Additionally, for comparison with InvRender [zhang2022modeling], we employ the 3D reconstruction from our Stage 1 and modify InvRender to incorporate object segmentation masks for computing pixel-wise losses in hand-object interaction images. We also adapt the approach by introducing Spherical Gaussian (SG) lobe rotations to effectively handle dynamic rigid object rotations, which are essential for handling both static and dynamic elements in the scene.

### v2

@@ 4.1.4 Pose Refinement
where $C(r)$ and $\hat{C}(r)$ are the predicted and groundtruth RGB colors along the ray $r$. Hand, object, and background masks are ensured consistent with ground truth by calculating the loss between predicted segmentation one-hot label ${M}(r)$, calculated using Eq. 9, and estimated ground truth segmentation obtained using SAM-Track [11] $\hat{M}(r)$, that takes image frames as input, and user input for target object in the image for the first frame, and outputs segmentation masks tracking the segmented object, following the same methodology as HOLD [17]:
$\mathcal{L}_{seg1}=\Sigma_{r}||{M}(r)-{\hat{M}}(r)||.$ (11)
@@
where $\lambda_{seg1}$ is set to $1.1$ and progressively decreased to $0.1$ until $30,000$ iterations, $\lambda_{eikonal1}$ is $1$, $\lambda_{hand-sdf1}$ is set to $5.0$ and $\lambda_{contact1}$ is initially set to 00 and progressively increased to $1.0$ until $30,000$ iterations.
@@
Figure 2: After Stage 1, for each (a) input image, predicted (b) object mask $M_{obj}$ and (c) hand-object mask $M_{ho}$ are calculated.
For a given frame (Fig. 2 (a)), this stage prepares the object’s implicit network, hand and object pose, object mask $M_{obj}$ (Fig. 2 (b)) and hand-object mask $M_{ho}$ (Fig. 2 (c)) for the second stage of our method.
@@ 4.2.1 Environment Illumination
$L_{d}(\omega_{i})=\Sigma_{j=1}^{128}G(\omega_{i};\psi_{j},\mu_{j},{\eta}_{j}).$ (14)
@@ 4.2.3 Hand Occlusion
One of the key challenges in rendering hand-object interactions is handling occlusion caused by the hand. To model this, we represent the hand as a set of $108$ parameterized spheres similar to power cells [50]. We first place all spheres $\{p_{i},r_{i}\}_{i=1}^{108}$ manually in the canonical MANO hand volume $\Psi$, with $p_{i}$ and $r_{i}$ being their centers and radii.
@@
Initially, the sphere centers $\{p_{i}\}_{i=1}^{108}$ are determined on the canonical hand skeleton, manually selected between the MANO joints. The sphere radii $\{r_{i}\}_{i=1}^{108}$ requires partitioning the surface of the hand model into individual spheres. These spheres define a power diagram that partitions the MANO hand volume $\Psi$ into a set of power cells. Each power cell $\Psi_{i}^{pow}$ consists of the points $x$ that are closest to a particular sphere $(c_{i},r_{i})$:
@@
where $d_{pow}(x,p_{i},r_{i})=||x-p_{i}||^{2}-r_{i}^{2}$ is the power distance between the point $x$ and the sphere $(p_{i},r_{i})$. The power diagram partitions all vertices $\{v_{j}\}$ of canonical MANO hand model into each individual set $\mathcal{V}_{i}=\{v_{j}|v_{j}\in\Psi_{i}^{pow}\}$. Whenever the MANO hand model is deformed by its pose and shape parameters, the center $\hat{p}_{i}$ and radius $\hat{r}_{i}$ of each sphere can be updated correspondingly by the following rule:
$\begin{split}\hat{p}_{i}&=\frac{1}{|\mathcal{V}_{i}|}\sum_{v_{j}\in\mathcal{V}_{i}}\hat{v}_{j},\\ \hat{r}_{i}&=\frac{1}{|\mathcal{V}_{i}|}\sum_{v_{j}\in\mathcal{V}_{i}}||\hat{n}_{j}\cdot(\hat{v}_{j}-\hat{p}_{i})||,\end{split}$ (18)
@@
This granular representation allows us to model how different parts of the hand obstruct the light and cast shadows on the object’s surface. For each surface point on the object, we divide its lobe of Spherical Gaussian (SG) into a grid of 64x64 patches, each representing a small region of the environment illumination and BRDF. Each one of the $108$ spheres representing MANO hand is projected onto the SG lobe, and the patches belonging to the occluded regions are collected. Finally, the fractional value of the integral of the occluded patches gives a measure of the extent to which each part of the hand blocks light from reaching the object. For the following integral of a generic Spherical Gaussian (ISG) in polar format [67] with lobe sharpness $\eta$,
@@
where $\phi^{\prime}$ is angle between the lobe direction of the Spherical Gaussian, and some direction $\omega_{i}$ along the hemisphere. $S(\theta,\phi,{\eta})$ represents the accumulated Spherical Gaussian over a patch of the hemisphere, where $\theta$ and $\phi$ are angular coordinates. This means when $\theta=\pi$ and $\phi=\pi$, it covers the whole hemisphere. To compute the fractional value for a given angular region, we normalize by dividing the integral result by $S(\pi,\pi,{\eta})$ [23], as shown below:
$\hat{S}(\theta,\phi,{\eta})=\frac{S(\theta,\phi,{\eta})}{S(\pi,\pi,{\eta})}.$ (20)
Iwasaki et. al. [23] call this normalized term as normalized ISG, and approximate it as follows:
$\hat{S}(\theta,\phi,\eta)\approx\frac{1}{1+e^{-g(\eta)(\theta-\pi/2)}}\frac{1}{1+e^{-h(\eta)(\phi-\pi/2)}},$ (21)
where they define $g$ and $h$ as fourth-degree polynomial function approximations over $\eta$ as follows:
$\begin{split}g(\eta)&=g_{4}(\eta/100)^{4}+g_{3}(\eta/100)^{3}+g_{2}(\eta/100)^{2}+g_{1}(\eta/100)\\ h(\eta)&=h_{4}(\eta/100)^{4}+h_{3}(\eta/100)^{3}+h_{2}(\eta/100)^{2}+h_{1}(\eta/100),\end{split}$ (22)
where $g_{4}$, $g_{3}$, $g_{2}$, $g_{1}$, $h_{4}$, $h_{3}$, $h_{2}$, $h_{1}$ are constant values $-2.6856e^{-6}$, $7e^{-4}$,$-0.0571$, $3.9529$, $17.6028$, $-2.6875e^{-6}$, $7e^{-4}$, $-0.0592$, $3.9900$, $17.5003$ respectively. The patches $\tilde{\Omega}$ occluded by hand spheres can be seen in Fig. 6. The normalized value helps us calculate the proportion of the spherical Gaussian that falls within a specific patch of the sphere. From Eq. 21 and Eq. 22, for a patch defined along some direction $\omega_{i}$ by the angular coordinates $(\theta_{0},\phi_{0})$, $(\theta_{1},\phi_{0})$, $(\theta_{0},\phi_{1})$, and $(\theta_{1},\phi_{1})$ (See Fig. 5), the final output is determined by computing the contributions from the patch corners:
$\begin{split}{\mathcal{F}(x,\theta_{0},\theta_{1},\phi_{0},\phi_{1})}=\hat{S}(\theta_{1},\phi_{1},{\eta})-\hat{S}(\theta_{1},\phi_{0},{\eta})\\ -\hat{S}(\theta_{0},\phi_{1},{\eta})+\hat{S}(\theta_{0},\phi_{0},{\eta}).\end{split}$ (23)
The sum over all the occluded patches gives us the fractional value $\mathcal{F}(x)$, and is the fraction of ISG over occluded patches $\tilde{\Omega}$. Therefore, we can approximate the integration of a partially-occluded Spherical Gaussian $G(\omega_{i})O(x,\omega_{i})$, with respect to the direction $\omega_{i}$ over the hemisphere $\Omega^{+}$ centered around the surface point $x$, using the above integral over the occluded patches $\tilde{\Omega}$ and its approximation $\mathcal{F}(x)$ as follows:
[…]
@@ 4.2.4 Final Illumination
We discussed direct illumination $L_{d}(\omega_{i})$ from environment, as a sum of Spherical Gaussian along direction $w_{i}$ of hemisphere in Eq. 14, and integral of SG with hand occlusion in Eq. 24. In the regions of hand occlusion, the hand casts its own indirect illumination, which can be perceived as the hand’s reflection on the object’s surface. Hence, we learn a constant indirect illumination $L_{i}$, starting from a generic skin color (#$\#E0AC69$) from an online search. The final illumination $L(\omega_{i},x)$ from Eq. 1 can thus be substituted as follows:
$L(\omega_{i},x)=L_{d}(\omega_{i})*(1-O(x,\omega_{i}))+L_{i}*O(x,\omega_{i}).$ (25)
@@
Combining illumination $L(\omega_{i},x)$ from Eq. (25), BRDF $\phi(x;\omega_{o},\omega_{i})$ from Eq. (15), and cosine term $\omega_{i}\cdot n$ from Eq. (26), we get the final color $c(\omega_{o},x)$ at the surface point $x$ with view direction $\omega_{o}$ as the integral equation given in PBR Eq. (1). To maintain the consistency of object shape, and avoid the hand’s involvement in object geometry and texture reconstruction, we propose the following loss functions. An RGB loss is calculated on the color $c^{ho}$ predicted on $N_{ho}$ sampled pixels in the intersection of the predicted object mask and groundtruth hand-object mask $M_{ho}$ (Fig. 2(c)):
@@
where $\hat{c}_{i}$ is groundtruth color of the $N_{ho}$ sampled pixels. The below mask loss calculates minimum SDF $S^{no}$ on the $N_{no}$ rays corresponding to the sampled pixels outside the intersection of the predicted object mask and ground truth object mask $M_{obj}$ (Fig. 2(b)):
@@
Just like in stage 1, $\mathcal{L}_{eikonal2}$ regularizes object’s geometry. The final loss optimized is as follows:
$\mathcal{L}_{s2}=\mathcal{L}_{rgb2}+\lambda_{mask2}\mathcal{L}_{mask2}+\lambda_{eikonal2}\mathcal{L}_{eikonal2},$ (29)
where $\lambda_{mask2}$ is set to $100.0$ and $\lambda_{eikonal2}$ is $0.1$.
@@ 4.3 Implementation Details
We utilize monocular camera capture datasets that encompass all viewpoints around the object, ensuring comprehensive coverage for high-fidelity texture and geometry learning. For the first stage, the 3D annotations estimated include hand-object pixel-wise segmentation [11, 28, 62, 61, 32, 19], hand MANO pose estimation [31], and object pose estimation [40, 41] in the camera-coordinate system.
The first stage of our pipeline involves the composite volumetric rendering of the hand, object, and background, following a pipeline similar to HOLD [17], using the initial estimates of the hand and object poses, along with segmentation masks. HOLD optimizes hand and object poses, geometry, and an initial estimate of texture.
Using the optimized poses and object geometry from Stage 1, we derive the object mask and hand-object segmentation mask. These are then used in the second stage, which focuses on physics-based surface rendering of the object, following a pipeline inspired by PhySG [67]. A key modification in our method is the computation of occlusion to account for hand-object interaction, based on the approach in Iwasaki et al. [23]. Prior works, such as PhySG, do not compute occlusion in this way, setting our method apart. In Stage 2, we optimize the following key parameters at a surface point $x$:
@@
Figure 7: Qualitative Comparison: TexHOI albedo $a(x)$ predictions avoid baked hand-occlusions and environment-illumination. Hand occlusions can be observed in “Hand-Object Interaction” $\mathcal{F}(x)$ (0-1 normalized). The final TexHOI RGB $c(\omega_{o},x)$ is calculated using Physics-Based Rendering, and is compared with other state-of-the-art approaches, HOLD [17] and InvRender [69].
@@
TABLE I: Quantitative comparison is performed between our method, TexHOI, and other state-of-the-art approaches that reconstruct geometry and texture. HOLD bakes hand-occlusions and illumination on its texture prediction, InvRender does not handle dynamic occlusions.
Figure 8: Ablation studies compare our model with the versions trained without indirect illumination, and without any occlusions. Illumination effects on surface albedo, and hand occlusions on surface can be observed. The scenes are in same order as Fig. 7
@@
TABLE II: Quantitative ablative comparison of our model with the versions without indirect illumination, and without hand occlusions. Without illumination, the interacting hand casts a darker texture. Without occlusion handling, artifacts occur in major hand grasp regions.
@@ 5.1 Dataset
  • HO3D: The training data of HO3D [21] dataset contains 9 objects from YCB [57] models. We use monocular frames of six different objects from the dataset. These six objects were specifically selected based on the presence of sufficient texture details, which is essential for evaluating the effectiveness of our texture reconstruction. The six objects (and their subjects) are - sugar box (“ShSu”), cracker box (“MC”), bleach cleanser (“ABF”), meat can (“GPMF”), mustard bottle (“SM”), and drilling machine (“MDF”). The $3$ left objects in training data - scissors, banana, and mug contain no detailed-texture information for comparison.
  • In-the-wild: We also use monocular frame sequences of three objects captured in real-world settings (“in-the-wild”) [17]. These objects provide further diversity in testing by presenting challenging hand-object interactions and varied environmental conditions. The objects are - bottle, kettle, and rubrics cube.
@@ 5.2 Baselines
We use HOLD [17] as our primary baseline for performance comparison due to its proven effectiveness in dynamic hand-object interaction tasks, where it outperforms methods such as DiffHOI [65] and BundleSDF [53]. This makes HOLD an appropriate benchmark for assessing the improvements introduced by our method. Our approach, however, goes beyond what HOLD achieves by focusing on texture reconstruction and illumination effects, areas that HOLD does not fully address. To highlight the improvements, we compare our full pipeline results with HOLD alone. This comparison demonstrates how our method enhances texture accuracy, particularly in capturing the complex interactions between hands and objects, such as shadows, reflections, and visibility changes, which are often overlooked in prior works. Additionally, for comparison with InvRender [69], we employ the 3D reconstruction from our Stage 1 and modify InvRender to incorporate object segmentation masks for computing pixel-wise losses in hand-object interaction images. We also adapt the approach by introducing Spherical Gaussian (SG) lobe rotations to effectively handle dynamic rigid object rotations, which are essential for handling both static and dynamic elements in the scene.